OpenStudy (anonymous):

2. Prove that the following sets are denumerable. a. 3N, the positive integer ultiples of 3. b. 3Z, the integer multiples of 3 c. {x: x element N and n<-12} d. {(x,y) element N x R: xy=1} 5. {x element Z: x=1(mod5)}

OpenStudy (freckles):

Cay you find a bijection between the naturals and 3*naturals?

OpenStudy (freckles):

natural numbers are {1,2,3,4,...} 3*natural numbers are {1(3)=3,2(3)=6,3(3)=9,3(4)=12,...}

OpenStudy (freckles):

Find a function between those two sets then prove the function is bijective. This one should be pretty obvious.